# Pinching-Antenna Systems (PASS): Architecture Designs, Opportunities, and Outlook

### v2

@@ III-B Multiplexing Architecture
  • Fully-connected Architecture: In this architecture, the signals from each RF chain are fed into all waveguides with the aid of power splitters. The fully-connected architecture addresses two key limitations of the sub-connected design. First, it enables each RF chain to connect to all PAs across all waveguides, thereby fully using the overall spatial DoFs to enhance communication performance. Second, the number of RF chains no longer needs to match the number of waveguides. This flexibility allows for either reducing the number of RF chains to save energy, as hybrid beamforing in millimeter-wave networks, or increasing it to further exploit the additional spatial DoFs provided by the massive PAs.
@@ IV-A PASS for Ubiquitous Indoor and Outdoor Coverage
The low-complexity structure of PASS enables it to be easily deployed on diverse structures in both outdoor and indoor scenarios, such as building facades, rooftops, roadside guardrails, and indoor ceilings or walls, as depicted in Fig. 4. PASS address the key challenge of conventional wireless communication, where communication links are often partially or totally blocked by obstacles and suffer from significant long-distance path loss due to those remotely deployed BSs. In particular, PASS establishes LoS-dominated and short-distance wireless channels with the widespread deployment of waveguides and the optimal activation of PAs near the terminals. By preventing wireless signals from being wasted or even interfering other users, PASS is capable of facilitating ubiquitous outdoor and indoor coverage in a cost-efficient and energy-efficient manner. Several attractive applications of PASS in indoor and outdoor scenarios will be discussed below.
@@ IV-B Interplay between PASS and NOMA
The combination of PASS and non-orthogonal multiple access (NOMA) provides a “win-win” solution for 6G and beyond. On the one hand, since multiple PAs on a specific waveguide can only be fed with the same set of signals, how to serve multiple users simultaneously becomes a critical challenge. To address this, it is natural to use NOMA to multiplex multiple users’ signal in the power domain over the same radio resource [9]. On the other hand, PASS also facilitates flexible NOMA operations. Recall the fact that NOMA relies on successive interference cancellation (SIC) to handle co-channel interference and therefore SIC order significantly affects the communication performance achieved by each NOMA user. Conventionally, a fixed distance-based SIC order is employed for NOMA, where users closer to the BS, so called strong users, are assigned higher SIC orders due to better channel conditions. In contrast, as shown in the bottom right of Fig. 4, PASS-based NOMA enables a more flexible SIC operation. This is achieved by dynamically adjusting users’ channel conditions through the allocation of PAs and the design of their activation locations. For example, users originally farther from the BS/AP can have higher channel conditions with more numbers of PAs allocated. By doing so, more customized wireless services can be provided by the PASS-based NOMA.
@@ IV-C PASS for NISE/ISAC
In addition to enhancing communication, PASS significantly improves sensing quality and coverage. Since the links between PAs and sensing targets follow the near-field model, PASS enables near-field sensing (NISE) to precisely estimate angular and range information, as well as target velocity [15]. As illustrated in the top left of Fig. 4, two practical structures can be employed: ISAC-PASS and NISE-PASS. (1) ISAC-PASS allocates part of the waveguides to transmit communication signals while using the rest to receive echo-sensing signals, enabling simultaneous near-field communication and sensing within the area of interest. (2) NISE-PASS, equipped exclusively with receiving waveguides, operates in a low-power mode to facilitate NISE. By leveraging the widespread deployment of PASS and successively or periodically activating sensing PAs, ubiquitous and highly accurate NISE can be achieved. This capability is particularly beneficial for real-world applications, such as tracking target trajectories or monitoring malicious entities.
Figure 6: Average transmit power versus the minimum SINR.
@@ V Numerical Case Studies
In this section, we present numerical examples to characterize the performance of PASS employing pinching beamforming and to compare them with the conventional baselines.
V-A Array Gains of PASS
The array gain of PASS is analyzed by considering a setup where multiple PAs, equally spaced along a single waveguide, are used to serve a single-antenna user; see [11, Figure 1]. This setup represents the most fundamental pinching beamforming scenario, and the corresponding results serve as a lower bound for those more advanced designs.
Unlike conventional fixed-location antenna systems, the wireless propagation in PASS is LoS-dominated and near-field-driven. Considering these factors, the array gain of the system has been analyzed with respect to the number of PAs, as illustrated in Figure 5 [11]. The findings in Figure 5 indicate that the array gain does not increase monotonically with the number of antennas. Instead, there exists an optimal number of PAs that maximizes the array gain achievable by PASS. Notably, the maximum array gain exceeds that of conventional-antenna systems. Since the considered naive pinching beamforming represents a lower bound for other potential designs, these observations lead to the following conclusions: i) Due to their flexibility, PASS can achieve a higher array gain than conventional fixed-location antenna systems. ii) Optimizing the number of PAs is essential for PASS to fully realize their performance potential.
V-B Transmit Power Minimization by PASS
Although PASS has demonstrated the ability to achieve enhanced array gains, the design of the joint transmit and pinching beamforming for maximizing the PASS performance is a non-trivial task. The key challenges stem not only from the coupling between transmit and pinching beamforming but also from the combined impact of all PAs on both the amplitude and phase of the signals. To tackle these challenges, we propose leveraging penalty dual decomposition (PDD) to effectively decompose the optimization parameters.
To evaluate the effectiveness of PDD in enhancing the performance of multi-user PASS, we examine a transmit power minimization problem constrained by the minimum signal-to-interference-plus-noise ratio (SINR) requirement. We assume a 15 GHz system where five 50-meter waveguides are deployed in parallel with a spacing of 5 meters, serving four users randomly distributed 15-20 meters away from the BS. Each waveguide has 6 PAs, and a sub-connected multiplexing transmission architecture is used. For comparison, we consider two baseline systems: fully-digital MIMO with 5 antennas and massive MIMO with 30 antennas and 5 RF chains. The numerical results obtained by the Monte Carlo method are illustrated in Fig. 6, where the number of available PA positions for discrete activation is set to 200. These results highlight the significant performance gains of PASS in reducing transmit power, even when compared to massive MIMO with a more complex transmission architecture. Additionally, it is interesting to observe that the performance loss caused by the proportional power model is negligible compared to the equal power model.
@@ VI Conclusions and Open challenges
In this article, the promising concept of PASS was introduced, highlighting their great potential for enabling efficient wireless communications over the “last meter”. The basic physical principle, signal model, and communication design of PASS have been discussed. By leveraging the novel pinching beamforming, two practical transmission architectures were developed for PASS, whose advantages and limitations were also described. Moreover, a range of promising applications for employing PASS in wireless networks were explored. Numerical case studies demonstrated the superiority of PASS over conventional multi-antenna transmission. As PASS remains in its early stages of research, there are numerous open challenges and opportunities to unlock its full potential. Several of these key research directions are exemplified below:
  • Spatial Modeling for PASS with Stochastic Geometry: Stochastic geometry provides a powerful framework for evaluating the average performance of wireless networks, which can be used to determine key design metrics, such as the required number of PAs, the ratio between PAs and users, and optimal placement strategies. However, the unique features of PASS—such as near-field effects, LoS-dominated links, and location-flexible pinching—necessitate the development of more accurate spatial models to capture the spatial randomness, which can be a promising and interesting research direction.
  • Machine Learning Empowered PASS: Machine learning empowers both channel estimation and fast pinching deployment for PASS. Owing to the large spatial extent, PASS may observe non-stationary multi-path characteristics in different serving areas. Generative learning can capture the spatial non-stationarity and reconstruct large-scale PASS channels. Furthermore, low-complexity learning-based designs are required for user-centric pinching of a large number of PAs, thus manipulating PAs for both path loss compensation and constructive interference.
  • Advanced Hardware Designs for PASS: Pinching beamforming plays a critical role in determining the flexibility of PASS and the maximum achievable performance gain. Advanced PASS hardware designs are essential in this context, posing several challenges, such as real-time and continuous control of PA movements or activation, active control of power leakage from PAs, mitigation of unwanted power losses, and desired dielectric materials. This requires multi-disciplinary research efforts.
@@
Yuanwei Liu (Fellow, IEEE) is a (tenured) Full Professor at The University of Hong Kong.
Zhaolin Wang (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.

### v3

@@ III-B Waveguide Multiplexing Architecture
  • Fully-connected Architecture: In this architecture, the signals from each RF chain are fed into all waveguides with the aid of power splitters. The fully-connected architecture addresses two key limitations of the sub-connected design. First, it enables each RF chain to connect to all PAs across all waveguides, thereby fully using the overall spatial DoFs to enhance communication performance. Second, the number of RF chains no longer needs to match the number of waveguides. This flexibility allows for either reducing the number of RF chains to save energy, as in hybrid beamforming for millimeter-wave networks, or increasing it to further exploit the additional spatial DoFs provided by the massive PAs.
@@ IV-A PASS for Ubiquitous Indoor and Outdoor Coverage
The low-complexity structure of PASS enables it to be easily deployed on diverse structures in both outdoor and indoor scenarios, such as building facades, rooftops, roadside guardrails, and indoor ceilings or walls, as depicted in Fig. 4. PASS addresses the key challenge of conventional wireless communication, where communication links are often partially or totally blocked by obstacles and suffer from significant long-distance path loss due to the remotely deployed BSs. In particular, PASS establishes LoS-dominated and short-distance wireless channels with the widespread deployment of waveguides and the optimal activation of PAs near the terminals. By preventing wireless signals from being wasted or even interfering with other users, PASS is capable of facilitating ubiquitous outdoor and indoor coverage in a cost-efficient and energy-efficient manner. Several attractive applications of PASS in indoor and outdoor scenarios will be discussed below.
@@ IV-B Interplay between PASS and NOMA
The combination of PASS and non-orthogonal multiple access (NOMA) provides a “win-win” solution for 6G and beyond. On the one hand, since multiple PAs on a specific waveguide can only be fed with the same set of signals, how to serve multiple users simultaneously becomes a critical challenge. To address this, it is natural to use NOMA to multiplex multiple users’ signals in the power domain over the same radio resource [9]. On the other hand, PASS also facilitates flexible NOMA operations. Recall the fact that NOMA relies on successive interference cancellation (SIC) to handle co-channel interference, and therefore SIC order significantly affects the communication performance achieved by each NOMA user. Conventionally, a fixed distance-based SIC order is employed for NOMA, where users closer to the BS, so-called strong users, are assigned higher SIC orders due to better channel conditions. In contrast, as shown in the bottom right of Fig. 4, PASS-based NOMA enables a more flexible SIC operation. This is achieved by dynamically adjusting users’ channel conditions through the allocation of PAs and the design of their activation locations. For example, users originally farther from the BS/AP can have higher channel conditions with more numbers of PAs allocated. By doing so, more customized wireless services can be provided by the PASS-based NOMA.
@@ IV-C PASS for NISE/ISAC
In addition to enhancing communication, PASS significantly improves sensing quality and coverage. Since the links between PAs and sensing targets follow the near-field model, PASS enables near-field sensing (NISE) to precisely estimate angular and range information, as well as target velocity. As illustrated in the top left of Fig. 4, two practical structures can be employed: ISAC-PASS and NISE-PASS. (1) ISAC-PASS allocates part of the waveguides to transmit communication signals while using the rest to receive echo-sensing signals, enabling simultaneous near-field communication and sensing within the area of interest. (2) NISE-PASS, equipped exclusively with receiving waveguides, operates in a low-power mode to facilitate NISE. By leveraging the widespread deployment of PASS and successively or periodically activating sensing PAs, ubiquitous and highly accurate NISE can be achieved. This capability is particularly beneficial for real-world applications, such as tracking target trajectories or monitoring malicious entities.
Figure 6: Minimum transmit power versus the minimum SINR. The detailed simulation setup is given in [12].
@@ V Numerical Case Studies
In this section, we present numerical examples to characterize the performance of PASS employing pinching beamforming and to compare it with the conventional baselines.
V-A Array Gains
The array gain of PASS is analyzed by considering a setup where multiple PAs, equally spaced along a single waveguide, are used to serve a single-antenna user; see [11, Figure 1]. This setup represents the most fundamental pinching beamforming scenario, and the corresponding results serve as a lower bound for those more advanced designs. Unlike conventional fixed-location antenna systems, the wireless propagation in PASS is LoS-dominated and near-field-driven. Considering these factors, the array gain of the system has been analyzed with respect to the number of PAs, as illustrated in Figure 5 [11]. The findings in Figure 5 indicate that the array gain of pinching antennas does not increase monotonically with the number of antennas. Instead, there exists an optimal number of PAs that maximizes the array gain achievable by PASS. Notably, the maximum array gain exceeds that of conventional antenna systems. This considered naive pinching beamforming represents a lower bound for other potential designs, whose observations lead to the following conclusions: i) Due to their flexibility, PASS can achieve a higher array gain than conventional fixed-location antenna systems. ii) Optimizing the number of PAs is essential for PASS to fully realize its performance potential.
V-B Signal Power Enhancement
In Fig. 6, we illustrate the effectiveness of PASS in enhancing signal power through the joint optimization of pinching beamforming and transmit beamforming. The penalty-based method is exploited to deal with this coupled optimization problem.Specifically, we consider a sub-connected architecture for PASS under a practical proportional power model. Both continuous activation and discrete activation are evaluated, where the latter incorporates candidate PAs to be pre-installed per half meter. For benchmarking, we employ two baseline schemes: a conventional MIMO system, where each RF chain connects to a single antenna, and a massive MIMO system, where each RF chain connects to multiple antennas via analog beamforming devices. All systems are configured with an identical number of RF chains, and only LoS channels are considered to enable a fair and equitable comparison. The results indicate that PASS significantly reduces the transmit power required to satisfy the signal-to-interference-plus-noise ratio (SINR) demands of communication users. Additionally, although discrete activation slightly limits beamforming flexibility compared to continuous activation, PASS still demonstrates substantial performance benefits. This improvement primarily stems from the establishment of robust LoS channels, effectively mitigating free-space path loss.
@@ VI Conclusions and Open challenges
In this article, the promising concept of PASS was introduced, highlighting its great potential for enabling efficient wireless communications over the “last meter”. The basic physical principle, signal model, and communication design of PASS have been discussed. By leveraging the novel pinching beamforming, two practical transmission architectures were developed for PASS, whose advantages and limitations were also described. Moreover, a range of promising applications for employing PASS in wireless networks were explored. Numerical case studies demonstrated the superiority of PASS over conventional multi-antenna transmission. As PASS remains in its early stages of research, there are numerous open challenges and opportunities to unlock its full potential. Several of these key research directions are exemplified below:
  • Spatial Modeling for PASS with Stochastic Geometry: Stochastic geometry provides a powerful framework for evaluating the average performance of wireless networks, which can be used to determine key design metrics, such as the required number of PAs, the ratio between PAs and users, and optimal placement strategies. However, the unique features of PASS—such as near-field effects, LoS-dominated links, and location-flexible pinching—necessitate the development of more accurate spatial models to capture the spatial randomness, which can be a promising and interesting research direction. Furthermore, although PASS system is LoS-dominated, realistic models covering LoS blockage and residual multipath are still essential for practical performance evaluation.
  • Machine Learning Empowered PASS: Machine learning empowers both channel estimation and efficient pinching deployment for PASS. Owing to the large spatial extent, PASS may observe non-stationary multi-path characteristics in different serving areas. Generative learning can capture the spatial non-stationarity and reconstruct accurate PASS channels. Moreover, low-complexity learning-based designs can handle the nonconvex and coupled optimization even for a large number of PAs, thus manipulating PAs for both path loss compensation and constructive interference.
  • Advanced Hardware Designs for PASS: Pinching beamforming plays a critical role in determining the flexibility of PASS and the maximum achievable performance gain. Advanced PASS hardware designs are essential in this context, posing several challenges, such as real-time and continuous control of PA movements or activation, active control of power leakage from PAs, mitigation of unwanted power losses, and desired dielectric materials. Furthermore, hardware intended for wideband transmission must also consider waveguide dispersion, cut-off frequency, and the frequency-dependent coupling factor. These require multi-disciplinary research efforts.
@@
Yuanwei Liu (Fellow, IEEE) is a Professor at The University of Hong Kong.
Zhaolin Wang (Member, IEEE) is a Research Assistant Professor at The University of Hong Kong.